# Challenge LLMs to Reason About Reasoning: A Benchmark to Unveil Cognitive Depth in LLMs

### 28 Dec 2023 v1

@@ Abstract
In this work, we introduce a novel evaluation paradigm for Large Language Models, one that challenges them to engage in meta-reasoning. This approach addresses critical shortcomings in existing math problem-solving benchmarks, traditionally used to evaluate the cognitive capabilities of agents. Our paradigm shifts the focus from result-oriented assessments, which often overlook the reasoning process, to a more holistic evaluation that effectively differentiates the cognitive capabilities among models. For example, in our benchmark, GPT-4 demonstrates a performance ten times more accurate than GPT3-5. The significance of this new paradigm lies in its ability to reveal potential cognitive deficiencies in LLMs that current benchmarks, such as GSM8K, fail to uncover due to their saturation and lack of effective differentiation among varying reasoning abilities. Our comprehensive analysis includes several state-of-the-art math models from both open-source and closed-source communities, uncovering fundamental deficiencies in their training and evaluation approaches. This paper not only advocates for a paradigm shift in the assessment of LLMs but also contributes to the ongoing discourse on the trajectory towards Artificial General Intelligence (AGI). By promoting the adoption of meta-reasoning evaluation methods similar to ours, we aim to facilitate a more accurate assessment of the true cognitive abilities of LLMs.
Challenge LLMs to Reason About Reasoning: A Benchmark to Unveil Cognitive Depth in LLMs
@@
Haiyun Jiang thanks: Corresponding author. Tencent AI Lab haiyunjiang@tencent.com                        Jiaya Jia thanks: Corresponding author. Chinese University of Hong Kong leojia@cse.cuhk.edu.hk
@@ 1 Introduction
Pretrained on trillions of tokens and possessed with billions of parameters, today’s large language model (OpenAI2023GPT4TR; ANTHROPIC2023CLAUDE; touvron2023llama) is capable of generating coherent texts that are grammatically correct and seemingly meaningful. If further enhanced with the instruction tuning and human values alignment (e.g. through reinforcement learning from human feedback ouyang2022training), their outputs are on par with and even more superior than that of human beings. With the hope of differentiating different model’s cognitive ability, math questions are often selected as a proxy evaluation task. However, despite the complexity and diversity of these math problems, recent SOTA LLMs (OpenAI2023GPT4TR; yu2023metamath; Gou2023ToRAAT) have been able to achieve accuracy rates exceeding 80% (luo2023wizardmath) on multi-step math reasoning datasets like GSM8K (cobbe2021training).
Upon a detailed examination of the design principles and objectives of current math datasets, we identified several key shortcomings. Firstly, the majority of these datasets focus on result-oriented metrics, scoring accuracy based solely on the final answer, without considering the underlying reasoning process. With the chain of thought methodology (Wei2022ChainOT) and its derivative techniques (Chen2022ProgramOT; Yao2023TreeOT) emerged as the de facto standard for reasoning processes, we argue that the result-driven evaluation method may be insufficient for a comprehensive assessment of the intended cognitive and reasoning capabilities. Secondly, as a recent study (PasterHungarianTesting) suggests that some LLMs who achieved SOTA performances in GSM8K and MATH (hendrycks2021measuring) benchmarks have unexpectedly low performance when facing newly released Hungarian high school exams. This raises concerns about the data contamination and potential overfitting to the benchmarks. It also challenges the efficacy of these benchmarks in differentiating model capabilities.
Figure 1: Comparative Performance in GSM8K and DiagGSM8K: This graph illustrates the performance of GPT-4, Claude2, and GPT3-5 in the standard GSM8K benchmark versus our novel DiagGSM8K benchmark. While these models show similar performance levels in GSM8K, a marked variance is observed in DiagGSM8K, highlighting the differentiating power of our benchmark in evaluating deeper reasoning capabilities.
@@
In response to these identified limitations, we have introduced a novel paradigm, namely challenging LLMs to reason about different reasoning. This approach is specifically designed to augment existing evaluation benchmarks (as illustrated in Figure-3). Under this paradigm, LLMs are tasked to adopt a role akin to that of a teacher, assessing solutions by determining correctness, identifying potential initial errors, and providing reasons for these errors. Following this design principle, we have developed a novel benchmark named DiagGSM8k. This benchmark, characterized by instances manually labeled by experts and rigorously reviewed, serves as a robust tool for a qualitative and quantitative assessment of language models. Our findings indicate that most SOTA models only manage to achieve single-digit accuracy in this more nuanced assessment. As demonstrated in Figure-1, although SOTA models exhibit comparable performance in GSM8K, there is a considerable variance in their effectiveness on our benchmark, with discrepancies up to tenfold.
We argue that our evaluation paradigm does not only provide a more differentiating metric that focuses on the reasoning process over mere computational outcomes, but also sheds light on fundamental deficiencies within the current evaluation and training approaches. As visualized by Figure-2, we observe that recent SOTA models—regardless of whether they are trained on datasets that augment the respective benchmarks in terms of difficulty (luo2023wizardmath), coverage (lee2023platypus; yue2023mammoth), reasoning types (yu2023metamath; yue2023mammoth), solution diversities (yuan2023scaling), solution types (Gou2023ToRAAT; yue2023mammoth), or through reinforcement learning (Uesato2022SolvingMW; luo2023wizardmath), are essentially trained by teaching the language models to output a few solution paths within a large search space. Although these models can generate seemingly correct solutions, their grasp of underlying rationale and principles is often superficial and unsophisticated. This is evidenced by issues such as unit inconsistency during calculations (Toh2023VerityMathAM) which highlights a lack of fundamental ontological understanding, and the inability to discern nuanced differences between various reasoning paths for the same problems it was trained on. These core cognitive understanding abilities, vital for benchmarks to assess, have been overlooked due to the prevalent result-oriented metrics and evaluation paradigms. For downstream applications, particularly in education and consulting, the critical factor for success lies in the capability to rigorously analyze different plausible solutions, offering a comprehensive and well-rounded exploration of the solution space. We will provide a more in-depth discussion in section-5.
@@
1. Introduction of a novel evaluation principle and the accompanying open-source benchmark DiagGSM8K.  [^1] We have open-sourced our benchmark and evaluation results in https://github.com/dvlab-research/DiagGSM8K This benchmark challenges language models to engage in meta-reasoning, adding a new dimension to the evaluation of cognitive abilities in LLMs, beyond traditional benchmarks.
2. Comprehensive analysis of several SOTA models using the DiagGSM8K benchmark, highlighting critical shortcomings in the current training and evaluation paradigms. This analysis provides valuable insights into the true cognitive capabilities of these models.
@@ 2 Related Works
Complex reasoning tasks like math problems have long been widely accepted as a great proxy to fathom the cognitive ability in language models (Sharples1989ComputersAT; KoncelKedziorski2016MAWPSAM; Szegedy2020APP; Polu2020GenerativeLM; Miao2020ADC; hendrycks2021measuring; cobbe2021training). It demands the ability to understand the symbols and texts behind the problems, to dissect the problems into logically connected sub-ones, then combine and arrange results into final solutions. It touches on the cognitive abilities to induce patterns out of problems, to recall corresponding formulae, apply the rules deductively and reason in abstract symbolic way.
@@
GSM8K (cobbe2021training) and MATH (hendrycks2021measuring) have been two popular benchmarks to evaluate the math reasoning capabilities of LLMs in the past few years. Wei2022ChainOT proposed chain of thought to approach the multi-step reasoning tasks in a step by step manner. cobbe2021training and Uesato2022SolvingMW suggested to train a outcome based verifier model and apply reinforcement learning to help improve the performance of language models on GSM8k. Along the reinforcement learning route, Lightman2023LetsVS and luo2023wizardmath applied process based verifier model to aid in the training and inference stage. Stanford Alpaca (alpaca) popularized the knowledge distillation method of cloning corresponding abilities from ChatGPT (OpenAI2022CHATGPT) by asking it to generate related QA pairs. WizardMath (luo2023wizardmath) enhanced the distillation by specifying the QA difficulties in the generation process. Mammoth (yue2023mammoth) combined chain of thought and program of thought, and finetunes its models with the answer generated by GPT-4 (OpenAI2023GPT4TR) that are either in natural language or code language. MetaMath (yu2023metamath) expanded the generated question types by introducing the forward/backward reasoning variations.
@@
Despite the remarkable progress made in math reasoning and other natural language understanding tasks such as MMLU (Hendrycks2020MeasuringMM) and BBH (Suzgun2022ChallengingBT), there are some evidence that shows large language models might not have really mastered the reasoning and even not understanding what it generated. For example, Dziri2023FaithAF evaluated a few compositional reasoning problems and found that LLMs fail to generalize to questions with different complexity than the questions the model was trained on. West2023TheGA found the paradox that AI models can confidently generate some output that surpass average human level but failed to present basic levels of understanding on the output generated. Arkoudas2023GPT4CR used a collection of 21 reasoning tasks to systematically demonstrated that, despite occasional flashes of analytical brilliance, GPT4 is still severely limited when it comes to reasoning. Huang2023LargeLM and Yen2023ThreeQC also found that ChatGPT is very limited in judging the solution correctness of math problems, however our work focuses on the construction of a qualitative and quantitative evaluation framework and focus on the discussion of the evaluation principle and deficiencies of current training paradigm.
@@
Figure 3: This figure illustrates the structure of the DiagGSM8K benchmark, which incorporates math problems from GSM8K cobbe2021training and introduces two additional variations: Program of Thought (POT) (yue2023mammoth) and backward reasoning (yu2023metamath). These variations serve as sophisticated proxy challenges to assess whether language models possess the ability to ’reason about reasoning.’ Each question-solution pair within the benchmark demands that the evaluated model not only determines the solution’s correctness but also identifies the first erroneous step and articulates the underlying error rationale. Note despite the simplicity of the problem, GPT4 fail to predict the correctness of the POT variation.
@@ 3 Evaluation Framework
As introduced in Section1, our hypothesis posits that language models, when trained with the aim of optimizing a limited set of reasoning paths within a vast solution space, may lack the capability to comprehend and diagnose alternative reasoning approaches. To test this hypothesis, we developed the DiagGSM8K benchmark, specifically designed to uncover deficiencies in the current training paradigms. As illustrated in Figure-3, given a GSM8K question and its solution, the evaluated model is tasked to predict the correctness of the solution. If the solution is incorrect, the model is expected to further locate the first error location and elucidate the error reason.
Note that each test problem is combined with two variations which requires code solution (yue2023mammoth) and backward reasoning (yu2023metamath). The variations types are chosen specifically due to their significance in expanding the reasoning methodologies in LLMs. The ’Program of Thought,’ a concept proposed by Madaan2022LanguageMO and empirically proven effective in math reasoning by yue2023mammoth; Gou2023ToRAAT, represents a robust reasoning framework. The reasoning process in code language is inherently more structured and hierarchically abstracted. Its reasoning process is also less prone to error types such as calculation error. Backward reasoning, on the other hand, is recently brought under spotlight with discussions on if language models are able to learn backward relations effectively (Berglund2023TheRC).
For each problem collected, we utilized MetaMath-7B (yu2023metamath) with a temperature setting of 1 to generate step by step solutions. A panel of selected annotators were then recruited to review each QA pair on its reasoning process and decide the solution correctness, first error step and error reason. Followings are the annotation procedures adopted by the annotators:
First Error Step: This attribute is applicable for solutions with either an unmatched final output or a matched final output underpinned by flawed reasoning. Annotators identify the initial step in the reasoning process where the logic deviates from correctness. In line with the approach of Lightman2023LetsVS, we dissected GSM8K solutions into discrete steps, each marked by a newline character, and indexed them accordingly. Each step is then categorized as positive, neutral, or negative. Positive and neutral steps represent stages in the reasoning process where the correct final output remains attainable, whereas negative steps indicate a divergence from the path leading to the correct solution.
@@
Table 1: This table presents the composition of the DiagGSM8K benchmark. It categorizes questions into ’Original’ (based on GSM8K), ’POT’ (Program of Thought), and ’Reverse’ types, with counts for both correct and incorrect reasoning processes. Due to the difficulties of labelling coding solutions, the POT types of problems are labelled by the author manually therefore smaller in size.
Table 1 shows the statistics about the evaluation benchmark we curated. For more details of the annotation process such as the process design, annotation format please refer to Appendix-LABEL:sec:appendix_annotation.

### 20 Jan 2024 v2

# MR-GSM8K: A Meta-Reasoning Revolution in Large Language Model Evaluation

@@ Abstract
In this work, we introduce a novel evaluation paradigm for Large Language Models, one that challenges them to engage in meta-reasoning. This approach addresses critical shortcomings in existing math problem-solving benchmarks, traditionally used to evaluate the cognitive capabilities of agents. Our paradigm shifts the focus from result-oriented assessments, which often overlook the reasoning process, to a more holistic evaluation that effectively differentiates the cognitive capabilities among models. For example, in our benchmark, GPT-4 demonstrates a performance five times better than GPT3-5. The significance of this new paradigm lies in its ability to reveal potential cognitive deficiencies in LLMs that current benchmarks, such as GSM8K, fail to uncover due to their saturation and lack of effective differentiation among varying reasoning abilities. Our comprehensive analysis includes several state-of-the-art math models from both open-source and closed-source communities, uncovering fundamental deficiencies in their training and evaluation approaches. This paper not only advocates for a paradigm shift in the assessment of LLMs but also contributes to the ongoing discourse on the trajectory towards Artificial General Intelligence (AGI). By promoting the adoption of meta-reasoning evaluation methods similar to ours, we aim to facilitate a more accurate assessment of the true cognitive abilities of LLMs.
MR-GSM8K: A Meta-Reasoning Revolution in Large Language Model Evaluation
@@
Shu Liu Smartmore Co.Ltd sliu@smartmore.com                        Haiyun Jiang thanks: Corresponding author. Tencent AI Lab haiyunjiang@tencent.com                        Jiaya Jia thanks: Corresponding author. Chinese University of Hong Kong leojia@cse.cuhk.edu.hk
@@ 1 Introduction
Pretrained on trillions of tokens and possessed with billions of parameters, today’s large language model (OpenAI, 2023; Anthropic, 2023; Touvron et al., 2023) is capable of generating coherent texts that are grammatically correct and seemingly meaningful. If further enhanced with the instruction tuning and human values alignment (e.g. through reinforcement learning from human feedback Ouyang et al., 2022), their outputs are on par with and even more superior than that of human beings. With the hope of differentiating different model’s cognitive ability, math questions are often selected as a proxy evaluation task. However, despite the complexity and diversity of these math problems, recent SOTA LLMs (OpenAI, 2023; Yu et al., 2023; Gou et al., 2023) have been able to achieve accuracy rates exceeding 80% (Luo et al., 2023) on multi-step math reasoning datasets like GSM8K (Cobbe et al., 2021).
Upon a detailed examination of the design principles and objectives of current math datasets, we identified several key shortcomings. Firstly, the majority of these datasets focus on result-oriented metrics, scoring accuracy based solely on the final answer, without considering the underlying reasoning process. With the chain of thought methodology (Wei et al., 2022) and its derivative techniques (Chen et al., 2022; Yao et al., 2023) emerged as the de facto standard for reasoning processes, we argue that the result-driven evaluation method may be insufficient for a comprehensive assessment of the intended cognitive and reasoning capabilities. Secondly, as a recent study (Paster, 2023) suggests that some LLMs who achieved SOTA performances in GSM8K and MATH (Hendrycks et al., 2021) benchmarks have unexpectedly low performance when facing newly released Hungarian high school exams. This raises concerns about the data contamination and potential overfitting to the benchmarks. It also challenges the efficacy of these benchmarks in differentiating model capabilities.
Figure 1: Comparative Performance in GSM8K and MR-GSM8k: This graph illustrates the performance of GPT-4, Claude2, and GPT3-5 in the standard GSM8K benchmark versus our novel MR-GSM8k benchmark. While these models show similar performance levels in GSM8K, a marked variance is observed in MR-GSM8k, highlighting the differentiating power of our benchmark in evaluating deeper reasoning capabilities.
@@
In response to these identified limitations, we have introduced a novel paradigm, namely challenging LLMs to reason about different reasoning. This approach is specifically designed to augment existing evaluation benchmarks (as illustrated in Figure-3). Under this paradigm, LLMs are tasked to adopt a role akin to that of a teacher, assessing solutions by determining correctness, identifying potential initial errors, and providing reasons for these errors. Following this design principle, we have developed a novel benchmark named Meta-Reasoning-GSM8k (e.g.MR-GSM8k). This benchmark, characterized by instances manually labeled by experts and rigorously reviewed, serves as a robust tool for a qualitative and quantitative assessment of language models. Our findings indicate that most SOTA models demonstrate a significant performance decline in this more nuanced assessment. As demonstrated in Figure-1, although SOTA models exhibit comparable performance in GSM8K, there is a considerable variance in their effectiveness on our benchmark, with discrepancies up to fivefold.
We argue that our evaluation paradigm does not only provide a more differentiating metric that focuses on the reasoning process over mere computational outcomes, but also sheds light on fundamental deficiencies within the current evaluation and training approaches. As visualized by Figure-2, we observe that recent SOTA models—regardless of whether they are trained on datasets that augment the respective benchmarks in terms of difficulty (Luo et al., 2023), coverage (Lee et al., 2023; Yue et al., 2023), reasoning types (Yu et al., 2023; Yue et al., 2023), solution diversities (Yuan et al., 2023), solution types (Gou et al., 2023; Yue et al., 2023), or through reinforcement learning (Uesato et al., 2022; Luo et al., 2023), are essentially trained by teaching the language models to output a few solution paths within a large search space. Although these models can generate seemingly correct solutions, their grasp of underlying rationale and principles is often superficial and unsophisticated. This is evidenced by issues such as unit inconsistency during calculations (Toh et al., 2023) which highlights a lack of fundamental ontological understanding, and the inability to discern nuanced differences between various reasoning paths for the same problems it was trained on. These core cognitive understanding abilities, vital for benchmarks to assess, have been overlooked due to the prevalent result-oriented metrics and evaluation paradigms. For downstream applications, particularly in education and consulting, the critical factor for success lies in the capability to rigorously analyze different plausible solutions, offering a comprehensive and well-rounded exploration of the solution space. We will provide a more in-depth discussion in Section-5.
@@
1. Introduction of a novel evaluation principle and the accompanying open-source benchmark MR-GSM8k. [^1] We have open-sourced our benchmark and evaluation results in https://github.com/dvlab-research/MR-GSM8KThis benchmark challenges language models to engage in meta-reasoning, adding a new dimension to the evaluation of cognitive abilities in LLMs, beyond traditional benchmarks.
2. Comprehensive analysis of several SOTA models using the MR-GSM8k benchmark, highlighting critical shortcomings in the current training and evaluation paradigms. This analysis provides valuable insights into the true cognitive capabilities of these models.
@@ 2 Related Works
Complex reasoning tasks like math problems have long been widely accepted as a great proxy to fathom the cognitive ability in language models (Sharples et al., 1989; Koncel-Kedziorski et al., 2016; Szegedy, 2020; Polu and Sutskever, 2020; Miao et al., 2020; Hendrycks et al., 2021; Cobbe et al., 2021). It demands the ability to understand the symbols and texts behind the problems, to dissect the problems into logically connected sub-ones, then combine and arrange results into final solutions. It touches on the cognitive abilities to induce patterns out of problems, to recall corresponding formulae, apply the rules deductively and reason in abstract symbolic way.
@@
GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) have been two popular benchmarks to evaluate the math reasoning capabilities of LLMs in the past few years. Wei et al., 2022 proposed chain of thought to approach the multi-step reasoning tasks in a step by step manner. Cobbe et al., 2021 and Uesato et al., 2022 suggested to train a outcome based verifier model and apply reinforcement learning to help improve the performance of language models on GSM8k. Along the reinforcement learning route, Lightman et al., 2023 and Luo et al., 2023 applied process based verifier model to aid in the training and inference stage. Stanford Alpaca (Taori et al., 2023) popularized the knowledge distillation method of cloning corresponding abilities from ChatGPT (OpenAI, 2022) by asking it to generate related QA pairs. WizardMath (Luo et al., 2023) enhanced the distillation by specifying the QA difficulties in the generation process. Mammoth (Yue et al., 2023) combined chain of thought and program of thought, and finetunes its models with the answer generated by GPT-4 (OpenAI, 2023) that are either in natural language or code language. MetaMath (Yu et al., 2023) expanded the generated question types by introducing the forward/backward reasoning variations.
@@
Despite the remarkable progress made in math reasoning and other natural language understanding tasks such as MMLU (Hendrycks et al., 2020) and BBH (Suzgun et al., 2022), there are some evidence that shows large language models might not have really mastered the reasoning and even not understanding what it generated. For example, Dziri et al., 2023 evaluated a few compositional reasoning problems and found that LLMs fail to generalize to questions with different complexity than the questions the model was trained on. West et al., 2023 found the paradox that AI models can confidently generate some output that surpass average human level but failed to present basic levels of understanding on the output generated. Arkoudas, 2023 used a collection of 21 reasoning tasks to systematically demonstrated that, despite occasional flashes of analytical brilliance, GPT4 is still severely limited when it comes to reasoning. Huang et al., 2023a and Yen and Hsu, 2023 also found that ChatGPT is very limited in judging the solution correctness of math problems, however our work focuses on the construction of a qualitative and quantitative evaluation framework and focus on the discussion of the evaluation principle and deficiencies of current training paradigm.
@@
Figure 3: This figure illustrates the structure of the MR-GSM8k benchmark, which incorporates math problems from GSM8K Cobbe et al. (2021) and introduces two additional variations: Program of Thought (POT) (Yue et al., 2023) and backward reasoning (Yu et al., 2023). These variations serve as sophisticated proxy challenges to assess whether language models possess the ability to ’reason about reasoning.’ Each question-solution pair within the benchmark demands that the evaluated model not only determines the solution’s correctness but also identifies the first erroneous step and articulates the underlying error rationale. Note despite the simplicity of the problem, GPT4 fail to predict the correctness of the POT variation.
@@ 3 Evaluation Framework
3.1 Dataset Construction
As introduced in Section1, our hypothesis posits that language models, when trained with the aim of optimizing a limited set of reasoning paths within a vast solution space, may lack the capability to comprehend and diagnose alternative reasoning approaches. To test this hypothesis, we developed the MR-GSM8k benchmark, specifically designed to uncover deficiencies in the current training paradigms. As illustrated in Figure-3, given a GSM8K question and its solution, the evaluated model is tasked to predict the correctness of the solution. If the solution is incorrect, the model is expected to further locate the first error location and elucidate the error reason.
Note that each test problem is combined with two variations which requires code solution (Yue et al., 2023) and backward reasoning (Yu et al., 2023). The variations types are chosen specifically due to their significance in expanding the reasoning methodologies in LLMs. The ’Program of Thought,’ a concept proposed by Madaan et al., 2022 and empirically proven effective in math reasoning by Yue et al., 2023; Gou et al., 2023, represents a robust reasoning framework. The reasoning process in code language is inherently more structured and hierarchically abstracted. Its reasoning process is also less prone to error types such as calculation error. Backward reasoning, on the other hand, is recently brought under spotlight with discussions on if language models are able to learn backward relations effectively (Berglund et al., 2023).
For each problem collected, we utilized MetaMath-7B (Yu et al., 2023) with a temperature setting of 1 to generate step by step solutions. A panel of selected annotators were then recruited to review each QA pair on its reasoning process and decide the solution correctness, first error step and error reason. Table 1 shows the statistics about the evaluation benchmark we curated. For more details of the annotation process such as the process design, annotation challenges please refer to Appendix-A.
@@ 3.1 Dataset Construction
Table 1: This table presents the composition of the MR-GSM8k benchmark. It categorizes questions into ’Original’ (based on GSM8K), ’POT’ (Program of Thought), and ’Reverse’ types, with counts for both correct and incorrect reasoning processes. Due to the difficulties of labelling coding solutions, the POT types of problems are labelled by the author manually therefore smaller in size.
3.2 Evaluation Metric
For each question-answer pair annotated, the evaluated model are supposed to decide the correctness of the solution and report the first error step and error reason if any. The solution correctness and first error step is scored automatically based on the manual annotation result. Only when the evaluated model correctly identified the incorrect solution and error step will its error reason be further examined manually or automatically by models. Therefore in order to provide a unified and normalized score to reflect the overall competence of the evaluated model, we hereby propose a novel metric named MR-Score.
MR-Score is a weighted combination of three metrics. The first one is the Matthews Correlation Coefficient (e.g. MCC) for the binary classification of solution correctness. The MCC score ranges from -1 to +1 with -1 means total disagreement between prediction and observation, 0 indicates near random performance and +1 represents perfect prediction. Here we interpret negative values as no better than random guess and set 0 as cut-off threshold for normalization purpose. The second metric is the ratio between numbers of solutions with correct first error step predicted and the total number of incorrect solutions. The third metrics is likewise the ratio between number of solutions with correct first error step plus correct error reason predicted and the total number of incorrect solutions. The formula of the MR-Score is then given by
$Score=w_{1}*\max(0,MCC)+w_{2}*ACC_{step}$ (1)
$\quad+w_{3}*ACC_{reason}$
where
$MCC=\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)\times(TP+FN)}}$ (2)
×$\quad\times\frac{1}{\sqrt{(TN+FP)\times(TN+FN)}}$
$\begin{split}\text{ACC}_{\text{step}}=\frac{NUM_{\text{correct\_first\_error\_step}}}{NUM_{\text{incorrect\_sols}}}\end{split}$ (3)
$\begin{split}\text{ACC}_{\text{reason}}=\frac{NUM_{\text{correct\_error\_reason}}}{NUM_{\text{incorrect\_sols}}}\end{split}$ (4)
For the weights $w_{1},w_{2}$ and $w_{3}$, they are chosen empirically to be 0.2, 0.3 and 0.5 by considering the difficulties of binary solution correctness classification task, multi-class first error step predictions and free-form error reason explanations. For extended discussion on the design of MR-Score, please check out Appendix-B.